# Stochastic heat flow by moments

### v1

@@ Abstract.
The critical two-dimensional Stochastic Heat Flow (SHF) is the scaling limit of the directed polymers in random environments and the noise-mollified Stochastic Heat Equation (SHE), at the critical dimension of two and near the critical temperature. The work [CSZ23a] proved that the discrete polymers converge to a universal (model-independent) limit, thereby identifying the limit as the SHF. In this work, we take a different approach: We formulate a set of axioms for the SHF, prove the uniqueness in law under these axioms, and also prove the existence by showing that the noise-mollified SHE converges to the SHF under this formulation. The set consists of three axioms commonly seen in stochastic flows and one axiom about the first four moments of the SHF over every time interval.
@@ 1. Introduction
The critical two-dimensional Stochastic Heat Flow (SHF) is an intricate yet universal object. It describes the scaling limit of the directed polymers in random environments. Introduced by [HH85], the model of directed polymers is one of the most studied instances of disordered systems; see [Com17] for a review. At the same time, the SHF describes the scaling limit of the noise-mollified Stochastic Heat Equation (SHE). The SHE is an important instance of stochastic PDEs, in part due to its connection to the Kardar–Parisi–Zhang (KPZ) equation [KPZ86], which has been intensively studied in one dimension; see [FS10, Cor12, Qua12, QS15, CW17, CS19] for reviews. The dimension of two offers great interest because that is exactly where the diffusive effect of the polymer or equation and the roughening effect of the noise balance. This phenomenon is sometimes called marginally relevant in disorder systems [CSZ17] and called critical in stochastic PDEs. Existing theories in stochastic PDEs [Hai13, Hai14, GJ14, GIP15, GP17, Kup16, GP18] apply to subcritical ones only. The SHF is further critical in the sense that the temperature is tuned around the critical value. Below the critical value, the scaling limit is Gaussian [CSZ17, CSZ20, Gu20, CCR23, NN23] but with a variance strictly larger than that of the underlying noise. Around the critical temperature, a rather intricate scaling limit, the SHF, arises.
@@
Much effort has been devoted to obtaining the SHF. The first indication that a non-trivial scaling limit, that is the SHF, should exist comes from [BC98]. Based on [AGHKH88, Chapter I.5] and [ABD95], the work showed the convergence of the second moment of the mollified SHE to an explicit limit. The explicit limit of the moment suggests an intricate limit of the process. In terms of chaos expansion, the limit of the moment accumulates the contribution of a divergent number of chaos. Convergence of higher moments are significantly more challenging, and was later obtained in [CSZ19b] for the third moment, and in [GQT21] for of all moments based on [Raj99, DR04]; see [Che22, Che24a, Che24b] for probabilistic studies of the moments. These results ensured the tightness of the model in question, but the uniqueness of the limit point and the universality of the limit remained open. This challenging and important open problem was solved by [CSZ23a]. Through a combination of a coarse-graining argument and Lindeberg’s exchange method, they showed that the finite-dimensional distributions (in time) of the discrete polymers are Cauchy and the limit is universal (model-independent), thereby identifying the limit as the SHF.
In this paper, we take a different, axiomatic approach. In Definition 1.2, we formulate four axioms for the SHF. The first three are common for stochastic flows. The fourth axiom, which is the important one, requires the first four moments of the SHF to be given by the explicit formula in [GQT21]. Our main result, Theorem 1.3, proves that these four axioms uniquely characterize the law of the SHF. To establish the existence under this formulation, we prove the convergence of the mollified SHE to the SHF in Proposition 1.4.
@@
Let us recall some notation. The parameter $\theta\in\mathbb{R}$ denotes the finite-tuning of the temperature (see (1.4)) and the SHF depends on $\theta$. The $n$th moment of the SHF are described by a semigroup of operators $\mathcal{Q}^{[\![n]\!],\theta}(t)$ on $L^{2}(\mathbb{R}^{2n})$, called the semigroup of the delta-Bose gas. This semigroup was constructed in [GQT21, Section 8] based on [Raj99, DR04]. We will recall the definition of $\mathcal{Q}^{[\![n]\!],\theta}(t)$ in Section 2.1. All functions in this paper are real-valued. Write $\langle f_{1},f_{2}\rangle:=\int_{\mathbb{R}^{d}}\mathrm{d}x\,f_{1}(x)f_{2}(x)$ for the inner product on $L^{2}(\mathbb{R}^{d})$, write $C_{\mathrm{c}}(\mathbb{R}^{d})$ for the space of continuous and compactly supported functions on $\mathbb{R}^{d}$, write $M_{+}(\mathbb{R}^{d})$ for the space of positive, regular Borel measures on $\mathbb{R}^{d}$ such that $\mu(K)<\infty$ for every compact $K$, write $\mu f:=\int_{\mathbb{R}^{d}}\mu(\mathrm{d}x)\,f(x)$, and equip $M_{+}(\mathbb{R}^{d})$ with the vague topology, namely the coarsest topology such that $\mu\mapsto\mu f$ is continuous for every $f\in C_{\mathrm{c}}(\mathbb{R}^{d})$. We need the product $\mu\bullet_{u}\nu$ similar to the one introduced in [CM24].
@@
Let $\star$ denote the convolution. Call $\{u_{\ell}\}_{\ell=1}^{\infty}\subset L^{2}(\mathbb{R}^{2})$ an $\boldsymbol{L^{2}}$-applicable mollifier if every $u_{\ell}$ is nonnegative, $\sup\{\|u_{\ell}\star\psi\|_{2}\,|\,\ell=1,2,\ldots,\|\psi\|_{2}\leq 1\}<\infty$, and, for every $\psi\in L^{2}(\mathbb{R}^{2})$, $\|u_{\ell}\star\psi-\psi\|_{2}\to 0$ as $\ell\to\infty$. For $\mu,\nu\in M_{+}(\mathbb{R}^{4})$ and nonnegative $u\in L^{2}(\mathbb{R}^{2})$, call $(\mu\bullet_{u}\nu)$ well-defined if, for every $f\in C_{\mathrm{c}}(\mathbb{R}^{4})$, the integral
@@
converges absolutely. In this case, by the Riesz–Markov–Kakutani representation theorem, see [Rud86, Theorems 2.14 and 2.17(b)], $(\mu\operatorname*{\bullet}_{u}\nu)\in M_{+}(\mathbb{R}^{4})$.
@@
    $Z=Z_{s,t}$ is an $M_{+}(\mathbb{R}^{4})$-valued, continuous process on $\mathbb{R}^{2}_{\scriptscriptstyle\leq}:=\{(s\leq t)\in\mathbb{R}^{2}\}$,
@@
    for all $s<t<u$, there exists an $L^{2}$-applicable mollifier $\{u_{\ell}\}_{\ell}$ (which can depend on $s,t,u$) such that $Z_{s,t}\bullet_{u_{\ell}}Z_{t,u}-Z_{s,u}$ is well-defined and converges vaguely in probability to 00 as $\ell\to\infty$,
@@
    the moment formula holds for $\overline{n}=4$: For $n=1,\ldots,\overline{n}$, $s<t$, and $g_{i},g^{\prime}_{i}\in L^{2}(\mathbb{R}^{2})$, we have that $\mathbf{E}[\prod_{i=1}^{n}(Z_{s,t}\ |g_{i}\otimes g^{\prime}_{i}|)]<\infty$ and that
    $\mathbf{E}\Big{[}\prod_{i=1}^{n}\big{(}Z_{s,t}\ g_{i}\otimes g^{\prime}_{i}\big{)}\Big{]}=\Big{\langle}\bigotimes_{i=1}^{n}g_{i},\mathcal{Q}^{[\![n]\!],\theta}(t-s)\bigotimes_{i=1}^{n}g^{\prime}_{i}\Big{\rangle}\ .$ (1.2)
@@
It may appear surprising that the moments suffice for uniquely characterizing the SHF. A common wisdom says that the moments grow very fast and likely result in an indeterminate moment problem. In one dimension, the $n$th moment of the SHE grows exponentially $n^{3}$ [Che15, CG20, DT21, GL23]. For the SHF, [Raj99] predicted a double exponential growth; see [GQT21, Remark 1.8]. The key, however, is to consider not just the one-point moments but the moments over all time intervals together. This way, the fast moment growth actually becomes irrelevant, and the first few moments suffice. Roughly speaking, the idea is to use the moments to obtain the time continuity of the flow, which, when combined with Lindeberg’s exchange method, gives the uniqueness. In Section 1.2, we use the Geometric Brownian Flow (GBF, defined there) as an elementary example to illustrate how this works. The GBM has the same one-point moments as the log-normal random variable, which has an indeterminate moment problem. Despite this, we show that the first four moments over all time intervals together uniquely characterize the GBF.
@@
Next we turn to the convergence of the noise-mollified SHE. Fix a nonnegative, compactly supported $\varphi\in C^{\infty}(\mathbb{R}^{2})$ with $\int_{\mathbb{R}^{2}}\mathrm{d}x\,\varphi=1$, let $\xi(t,x)$ denote the spacetime white noise on ×$\mathbb{R}\times\mathbb{R}^{2}$, mollify the noise in space to get $\xi_{\varepsilon}(t,x):=\int_{\mathbb{R}^{2}}\mathrm{d}x^{\prime}\varphi((x-x^{\prime})\varepsilon^{-1})\varepsilon^{-2}\xi(t,x^{\prime})$, let
@@
$\partial_{t}\mathsf{Z}^{\varepsilon}=\tfrac{1}{2}\Delta\mathsf{Z}^{\varepsilon}+\sqrt{\beta_{\varepsilon}}\,\xi_{\varepsilon}\,\mathsf{Z}^{\varepsilon}\ .$ (1.5)
Let $\mathsf{Z}^{\varepsilon}_{s,t}(x^{\prime},x)$ be the fundamental solution of (1.5). Namely, $\mathsf{Z}^{\varepsilon}_{s,\cdot}(x^{\prime},\cdot)$ solves (1.5) for $(t,x)\in(s,\infty)\times\mathbb{R}^{2}$ with the initial condition $\mathsf{Z}^{\varepsilon}_{s,s}(x^{\prime},\cdot)=\delta_{x^{\prime}}$. By [Dal99], the solution uniquely exists and is measurable with respect to $\sigma(\{\xi(u,x)|s\leq u\leq t,x\in\mathbb{R}^{2}\})$. Further, by [CH19, Corollary 1.4], $\mathbf{P}[\mathsf{Z}^{\varepsilon}_{s,t}(x,x^{\prime})\geq 0,\forall(s,t,x,x^{\prime})\in\mathbb{R}^{2}_{\scriptscriptstyle\leq}\times\mathbb{R}^{2}]=1$. Set $Z^{\varepsilon}_{s,t}:=\mathrm{d}x^{\prime}\otimes\mathrm{d}x\,\mathsf{Z}^{\varepsilon}_{s,t}(x^{\prime},x)\in M_{+}(\mathbb{R}^{4})$. Equip $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},M_{+}(\mathbb{R}^{4}))$ with the uniform-on-compact topology. More explicitly, in Section 2.2, we give a metric $d$ that mertizes the vague topology on $M_{+}(\mathbb{R}^{4})$, which then induces the uniform-on-compact topology through
$\rho(\mu,\nu):=\sum_{\ell=1}^{\infty}2^{-\ell}\wedge\sup_{s\leq t\in[-\ell,\ell]}d(\mu_{s,t},\nu_{s,t})\ ,\qquad\mu,\nu\in C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},M_{+}(\mathbb{R}^{4}))\ .$ (1.6)
@@
Under the uniform-on-compact topology, the process $Z^{\varepsilon}$ converges in law to the SHF($\theta$). Hence the SHF($\theta$) exists (and is unique by Theorem 1.3).
1.1. Corollaries and future directions
As an immediate corollary of Theorem 1.3, the SHF($\theta$) enjoys a few invariants. The finite-dimensional (in time) analogs of these invariants were obtained in [CSZ23a, Theorem 1.2]. Let $(\tau_{z}g)(x):=g(x-z)$, with $x,z\in\mathbb{R}^{2}$, denote the shift operator on $\mathbb{R}^{2}$, and, for $r>0$, let $(\sigma^{n}_{r}f)(x):=f(x/r)$ denote the scaling operators on $\mathbb{R}^{2n}$.

### v3

@@ Abstract.
The Stochastic Heat Flow (SHF) emerges as the scaling limit of directed polymers in random environments and the noise-mollified Stochastic Heat Equation (SHE), specifically at the critical dimension of two and near the critical temperature. The prior work [CSZ23a] established the first construction of finite-dimensional distributions by demonstrating the universal (model-independent) convergence of discrete polymers. In this work, we present a new, independent approach to the SHF. We formulate the SHF as a continuous process and provide a set of axioms for its characterization. We establish both the uniqueness and existence of this process under our new formulation, with a key feature of these axioms being the matching of the first four moments.
Key words and phrases:
Axiomatic characterization, moments, stochastic heat flow
2020 Mathematics Subject Classification:
82C27, 60H17
@@ 1. Introduction
The critical two-dimensional Stochastic Heat Flow (SHF) is an intricate yet universal object. It describes the scaling limit of the directed polymers in random environments. Introduced by [HH85], the model of directed polymers is one of the most studied instances of disordered systems; see [Com17] for a review. At the same time, the SHF describes the scaling limit of the noise-mollified Stochastic Heat Equation (SHE). The SHE is an important instance of stochastic PDEs, in part due to its connection to the Kardar–Parisi–Zhang (KPZ) equation [KPZ86], which has been intensively studied in one dimension; see [FS10, Cor12, Qua12, QS15, CW17, CS19] for reviews. The dimension of two offers great interest because that is exactly where the diffusive effect of the polymer or equation and the roughening effect of the noise balance. This phenomenon is sometimes called marginally relevant in disorder systems [CSZ17] and called critical in stochastic PDEs. Most existing theories of singular stochastic PDEs [Hai13, Hai14, GJ14, GIP15, GP17, Kup16, GP18] require the equation of interest to be subcritical. The SHF is further critical in the sense that the temperature is tuned around the critical value. Below the critical value, the scaling limit is Gaussian [CSZ17, CSZ20, Gu20, CCR23, NN23] but with a variance strictly larger than that of the underlying noise. Around the critical temperature, a rather intricate scaling limit, the SHF, arises.
@@
Much effort has been devoted to obtaining the SHF. The first indication that a non-trivial scaling limit, that is the SHF, should exist comes from [BC98]. Based on [AGHKH88, Chapter I.5] and [ABD95], the work showed the convergence of the second moment of the mollified SHE to an explicit limit. The explicit limit of the moment suggests an intricate limit of the process. Convergence of higher moments are significantly more challenging, and was later obtained in [CSZ19b] for the third moment, and in [GQT21] for all moments based on [Raj99, DR04]; see [Che22, Che24a, Che24b] for probabilistic studies of the moments. These results ensured the tightness of the model in question, but the uniqueness of the limit point and the universality of the limit remained open. This challenging and important open problem was solved by [CSZ23a]. Through a combination of a coarse-graining argument and Lindeberg’s exchange method, they showed that the finite-dimensional distributions (in time) of the discrete polymers are Cauchy and the limit is universal (model-independent).
Still, it begged the question of describing the SHF without reference to pre-limit models and the question of constructing the full process beyond the finite-dimensional distributions.
In this paper, we take a new, independent approach and answer those questions. In Definition 1.2, we formulate the SHF as a continuous process and give a few axioms to characterize it. Our main result, Theorem 1.3, proves that these axioms uniquely characterize the law of the SHF, thereby offering a description of the SHF without reference to pre-limiting models. Next, based on this new axiomatic formulation, we establish the convergence of the mollified SHE to the SHF as a continuous process in Proposition 1.5, thereby giving the construction of the full process.
@@
Stating the main result requires some notation. The parameter $\theta\in\mathbb{R}$ denotes the fine-tuning of the temperature (see (1.6)) and the SHF depends on $\theta$. The $n$th moment of the SHF are described by a semigroup of operators $\mathcal{Q}^{[\![n]\!],\theta}(t)$ on $L^{2}(\mathbb{R}^{2n})$, called the semigroup of the delta-Bose gas. This semigroup was constructed in [GQT21, Section 8] based on [Raj99, DR04]. We will recall the definition of $\mathcal{Q}^{[\![n]\!],\theta}(t)$ in Section 2.1. Write $\langle f_{1},f_{2}\rangle:=\int_{\mathbb{R}^{d}}\mathrm{d}x\,f_{1}(x)f_{2}(x)$ for the inner product on $L^{2}(\mathbb{R}^{d})$, write $C_{\mathrm{c}}(\mathbb{R}^{d})$ for the space of continuous and compactly supported functions on $\mathbb{R}^{d}$, write $\mathcal{M}_{+}(\mathbb{R}^{d})$ for the space of locally finite (finite on compact sets), positive Borel measures on $\mathbb{R}^{d}$, write $\mu f:=\int_{\mathbb{R}^{d}}\mu(\mathrm{d}x)\,f(x)$, and equip $\mathcal{M}_{+}(\mathbb{R}^{d})$ with the vague topology, namely the coarsest topology such that $\mu\mapsto\mu f$ is continuous for every $f\in C_{\mathrm{c}}(\mathbb{R}^{d})$. We need the product $\mu\bullet_{u}\nu$ similar to the one introduced in [CM24].
@@
Let $\star$ denote the convolution. Call $\{u_{\ell}\}_{\ell=1}^{\infty}\subset L^{2}(\mathbb{R}^{2})\cap C(\mathbb{R}^{2})$ an $\boldsymbol{L^{2}}$-applicable mollifier if every $u_{\ell}$ is nonnegative and if $\phi\mapsto u\star\phi$ defines a bounded operator on $L^{2}(\mathbb{R}^{2})$ that strongly converges to the identity operator as $\ell\to\infty$. For $\mu,\nu\in\mathcal{M}_{+}(\mathbb{R}^{4})$ and nonnegative $u\in L^{2}(\mathbb{R}^{2})\cap C_{\mathrm{c}}(\mathbb{R}^{2})$, call $(\mu\bullet_{u}\nu)$ well-defined if, for every $f\in C_{\mathrm{c}}(\mathbb{R}^{4})$, the integral
@@
converges absolutely. In this case, by the Riesz–Markov–Kakutani representation theorem, see [Rud86, Theorems 2.14], $(\mu\operatorname*{\bullet}_{u}\nu)\in\mathcal{M}_{+}(\mathbb{R}^{4})$.
@@
    $Z=Z_{s,t}$ is an $\mathcal{M}_{+}(\mathbb{R}^{4})$-valued, continuous process on $\mathbb{R}^{2}_{\scriptscriptstyle\leq}:=\{(s\leq t)\in\mathbb{R}^{2}\}$,
@@
    for all $t_{0}<t_{1}<t_{2}$, there exists an $L^{2}$-applicable mollifier $\{u_{\ell}\}_{\ell}$ (which can depend on $t_{0},t_{1},t_{2}$) such that $Z_{t_{0},t_{1}}\bullet_{u_{\ell}}Z_{t_{1},t_{2}}-Z_{t_{0},t_{2}}$ is well-defined and converges vaguely in probability to 0 as $\ell\to\infty$,
@@
    the moment formula holds up to $\overline{n}=4$: For $n=1,\ldots,\overline{n}$, $s<t$, and $g_{i},g^{\prime}_{i}\in L^{2}(\mathbb{R}^{2})$,
    $\mathbf{E}\Big[\prod_{i=1}^{n}\,Z_{s,t}\ g_{i}\otimes g^{\prime}_{i}\Big]=\Big\langle\bigotimes_{i=1}^{n}g_{i},\mathcal{Q}^{[\![n]\!],\theta}(t-s)\bigotimes_{i=1}^{n}g^{\prime}_{i}\Big\rangle\ .$ (1.2)
Here is the main result of this paper, which concerns the uniqueness of SHF($\theta$). As said, the existence follows from Proposition 1.5 below.
@@
This new approach has some consequences. First, it allows a more straightforward proof of the convergence to the SHF, avoiding the coarse-graining procedure in [CSZ23a]. This improvement comes from a qualitative difference between the present work and the latter work, particularly in how Lindeberg’s method is used; see the third paragraph in Section 1.1 for more explanation. We demonstrate this point by proving the convergence of the noise-mollified SHE in Proposition 1.5. Next, and more importantly, the axiomatic characterization facilitates the study of the SHF. One example is the recent work [CT25] that proves the conditional Gaussian Multiplicative Chaos (GMC) structure of the associated polymer measure. The proof relies on the axiomatic characterization (Theorem 1.9 to be precise), and, in turn, the result yields further properties of the SHF; see Section 1.3 for more discussions. Another example is the recent work [GT25]. Based on the axiomatic formulation, the work showed that the SHF is a black noise in the sense of [TV98, Tsi04]. A crucial (yet subtle) fact used there is that the axiomatic formulation does not have a direct reference to pre-limiting models.
The SHF is one of the few instances where the limit of a critical stochastic PDE can be obtained. What makes critical equations difficult is that the driving noise typically becomes independent of the process in the limit. Let $Z^{\theta,\varepsilon}_{s,t}$ be the solution of the 2dd mollified SHE, which we define in Section 1.2. For $\varepsilon>0$, $Z^{\theta,\varepsilon}$ is measurable with respect to the spacetime white noise $\xi$, but as $\varepsilon\to 0$, $(Z^{\theta,\varepsilon},\xi)$ converges in law to ((SHF($\theta$$),\xi$), where SHF($\theta$) and $\xi$ are independent; see [GT25, Corollary 1.2] and [CD25]. This property rules out pathwise descriptions of the limiting process like those for subcritical equations. In turn, the distributional characterization Definition 1.2 seems natural and fitting.
We mention two more instances of critical-stochastic-PDE limits: the anisotropic KPZ (aKPZ) equation and Stochastic Burgers Equation (SBE). Under a weak scaling that logarithmically attenuates the nonlinearity, [CET23a, CES21] obtained the tightness and Gaussian limit of the 2dd aKPZ equation. Later, [CET20, CET23b] showed that, without the weak scaling, the aKPZ equation exhibits a surprising logarithmic superdiffusive behavior. Under a similar weak scaling, [CGT24] obtained the Gaussian limit for the SBE in 2dd and above. Later, for the 2dd SBE without the weak scaling, [DGHS24] obtained Tauberian-sense variance bounds, and [CMT24] obtained the Gaussian limit of the full process.
1.1. Moments, Lindeberg’s method, proof ideas
@@
It may appear surprising that the moments suffice for uniquely characterizing the SHF. Common wisdom says that the moments grow very fast and likely cause an indeterminate moment problem. In one dimension, the $n$th moment of the SHE grows exponentially $n^{3}$ [Kar87, Che15, LDMS16, CG20, DT21, GL23]. For the SHF, [Raj99] predicted a double exponential growth; see [GQT21, Remark 1.8]. Hence, for fixed $s<t$, the moment problem of $Z_{s,t}$ is likely ill-posed. The key, however, is to consider not just the one-point moments but the moments over all time intervals together. As explained below, doing so suffices for uniquely characterizing the SHF.
The first step in our proof is to use Lindeberg’s method to reduce the task to bounding moments. This step is carried out in Section 3.1. Let us explain the idea. Taking $Z$ and $\widetilde{Z}$ that satisfy Definition 1.2, we seek to show that $Z_{0,1}=\widetilde{Z}_{0,1}$ in law. As we will see in Section 2.2, Definition 1.2(d) allows us to extend the Chapman–Kolmogorov equation in Definition 1.2(b) to multifold, so
$Z_{0,1}=Z_{0,\frac{1}{N}}\operatorname*{\bullet}\cdots\operatorname*{\bullet}Z_{\frac{N-1}{N},1}\ .$ (1.3)
[…]
$Z_{0,1}=\big(\mathbf{E}Z_{0,\frac{1}{N}}+W_{0,\frac{1}{N}}\big)\operatorname*{\bullet}\cdots\operatorname*{\bullet}\big(\mathbf{E}Z_{\frac{N-1}{N},1}+W_{\frac{N-1}{N},1}\big)\ .$ (1.4)
Indeed, $\mathbf{E}Z=\mathbf{E}\widetilde{Z}$ by Definition 1.2(d), so the error comes from replacing $W$ with $\widetilde{W}$. As is common in using Lindeberg’s method, if we require the first $(\overline{n}-1)$ moments of $W$ and $\widetilde{W}$ to match, we can bound the error by the $p$th absolute moments for $p>\overline{n}-1$.
We emphasize that our way of using Lindeberg’s method qualitatively differs from that of [CSZ23a]. In the latter paper, Lindeberg’s method is used in the coarse-graining procedure through the chaos expansion of the pre-limiting model. The chaos expansion exhibits a singular behavior where the main contribution of the chaos “escapes” to ever-higher order in the scaling limit, a manifestation of the critical nature of the SHF. In the present paper, as explained above, Lindeberg’s method is used at the level of Definition 1.2 without referencing or using the chaos expansion.
Given the first step, the task boils down to finding an absolute $p$th moment of $W_{s,t}$ that decays faster than |ts||t-s| as $t-s\to 0$. Careful calculations show that the 2nd moment fails, which is unsurprising because that is also the case for the Brownian motion. At the same time, the calculations suggest that any $p>2$ should suffice. We take $\overline{n}=p=4$ because that is the smallest even power (even because we want absolute moments) above 2, and we only have explicit access to positive-integer moments. In Section 3.2, we bound the positive even moments of $W_{s,t}$ by utilizing the explicit expression of the delta-Bose semigroup $\mathcal{Q}^{[\![n]\!],\theta}$.
Remark 1.4.
One could seek to weaken Definition 1.2(d) to be a bound on the absolute $p$th moment of $W:=Z-\mathbf{E}Z$ for one $p>2$ and the matching of the $n=1,2$ moments. We do not do this because our approach to the Chapman–Kolmogorov equation (in Section 2.2) still requires $\overline{n}\geq 4$. This reliance can be removed by using the martingale approach by [CM24], with the specific choice of the $L^{2}$-applicable mollifier (Definition 1.1) being the heat kernel.
1.2. Convergence
Theorem 1.3 reduces proving the convergence to the SHF to controlling moments. Several pre-limiting models satisfy certain analogs of Definition 1.2(b)(c). For such models, proving the convergence boils down to proving tightness in a topology compatible with Definition 1.2(a) and verifying Definition 1.2(d). Indeed, tightness can be obtained by bounding moments, so the task comes down to bounding and showing the convergence of the moments.
@@ 1.2. Convergence
We demonstrate this point by showing the convergence of the noise-mollified SHE. To state the result, fix a nonnegative, compactly supported $\varphi\in C^{\infty}(\mathbb{R}^{2})$ with $\int_{\mathbb{R}^{2}}\mathrm{d}x\,\varphi=1$, let $\xi(t,x)$ denote the spacetime white noise on $\mathbb{R}\times\mathbb{R}^{2}$, mollify the noise in space to get $\xi_{\varepsilon}(t,x):=\int_{\mathbb{R}^{2}}\mathrm{d}x^{\prime}\varphi((x-x^{\prime})\varepsilon^{-1})\varepsilon^{-2}\xi(t,x^{\prime})$, let
@@
$\partial_{t}\mathsf{Z}^{\theta,\varepsilon}=\tfrac{1}{2}\Delta\mathsf{Z}^{\theta,\varepsilon}+\sqrt{\beta_{\varepsilon}}\,\xi_{\varepsilon}\,\mathsf{Z}^{\theta,\varepsilon}\ .$ (1.7)
Let $\mathsf{Z}^{\theta,\varepsilon}_{s,t}(x^{\prime},x)$ be the fundamental solution of (1.7). Namely, $\mathsf{Z}^{\theta,\varepsilon}_{s,\cdot}(x^{\prime},\cdot)$ solves (1.7) for $(t,x)\in(s,\infty)\times\mathbb{R}^{2}$ with the initial condition $\mathsf{Z}^{\theta,\varepsilon}_{s,s}(x^{\prime},\cdot)=\delta_{x^{\prime}}$. By [Dal99], the solution uniquely exists and is measurable with respect to $\sigma(\xi(u,x)|(u,x)\in[s,t]\times\mathbb{R}^{2})$. Further, by [CH19, Corollary 1.4], $\mathbf{P}[\mathsf{Z}^{\theta,\varepsilon}_{s,t}(x,x^{\prime})\geq 0,\forall(s,t,x,x^{\prime})\in\mathbb{R}^{2}_{\scriptscriptstyle\leq}\times\mathbb{R}^{2}]=1$. Set
$Z^{\theta,\varepsilon}_{s,t}:=\mathrm{d}x^{\prime}\otimes\mathrm{d}x\,\mathsf{Z}^{\theta,\varepsilon}_{s,t}(x^{\prime},x)\in\mathcal{M}_{+}(\mathbb{R}^{4})\ .$ (1.8)
Equip $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathcal{M}_{+}(\mathbb{R}^{4}))$ with the uniform-on-compact topology. More explicitly, in Section 2.2, we give a metric $d$ that metrizes the vague topology on $\mathcal{M}_{+}(\mathbb{R}^{4})$, which then induces the uniform-on-compact topology through
$\rho(\mu,\nu):=\sum_{\ell=1}^{\infty}2^{-\ell}\wedge\sup_{s\leq t\in[-\ell,\ell]}d(\mu_{s,t},\nu_{s,t})\ ,\qquad\mu,\nu\in C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathcal{M}_{+}(\mathbb{R}^{4}))\ .$ (1.9)
The following proposition will be proven in Section 4.
@@
Under the uniform-on-compact topology, $Z^{\theta,\varepsilon}$ converges in law to SHF($\theta$).
1.3. Polymer measure, conditional GMC, noise
The axiomatic characterization has been used in the studied properties of the SHF and the associate polymer measure. First, the polymer measure $M^{\theta}=M^{\theta}_{[s,t]}$ was recently constructed by [CM24]. The overall idea follows the construction in $1+1$ dimensions by [AKQ14], but the required Chapman–Kolmogorov equation becomes much more challenging in $1+2$ dimensions. The work [CM24] overcame this challenge by devising a martingale approach to the Chapman–Kolmogorov equation. In Section 2, we carry out a different approach via moments, which allows more general mollifiers as in Definition 1.1. Given the polymer measure, an interesting question asks whether it is a GMC. Indeed, for $\varepsilon>0$, one can check that $\mathsf{Z}^{\theta,\varepsilon}_{s,t}$ induces a polymer measure that is a GMC in the path space $C([s,t],\mathbb{R}^{2})$. The question asks whether the path-space GMC survives the $\varepsilon\to 0$ limit. This is the case in $1+1$ dimensions, as shown in [QRV22]. In $1+2$ dimensions, however, the work [CM24] explained in their Section 2.6 that $M^{\theta}$ cannot be a path-space GMC, and the work [CSZ23b] showed that the endpoint distribution of $M^{\theta}$ is not a GMC. On the other hand, Clark conjectured a conditional GMC property, based on previous works on polymers on the diamond hierarchical lattice [Cla23] and [ACK17, AC19, Cla19, Cla21, Cla22, CL23]. The conjecture states that, if one takes $M^{\theta}$, which is random, as the reference measure, then conditioned on $M^{\theta}$ and for $\theta^{\prime}>\theta$, the measure $M^{\theta^{\prime}}$ can be realized as a GMC with noise strength $\theta^{\prime}-\theta$.
The conditional GMC structure has recently been proven by [CT25] and has been used to derive a few properties, including the strict positivity of the SHF tested against general test functions. A key input to the proof is the axiomatic characterization from the present paper, more precisely Theorem 1.9. The characterization allows one to show, via moment calculations, that the conditional GMC with the reference measure $M^{\theta}$ and noise strength $\theta^{\prime}-\theta$ has the same law as $M^{\theta^{\prime}}$.
A potential future direction is to study the SHF as a noise in the sense of [TV98, Tsi04]. Based on the axiomatic formulation from the present paper, the recent work [GT25] proved that the SHF is a black noise, again in the sense of [TV98, Tsi04]. On a related note, the work [HP24] proved that the directed landscape, which the large-time limit of the logarithm of the one-dimensional SHE [Wu23] (see also [QS23, Vir20]), is a black noise.
1.4. Corollaries, generalizations
Let us mention a few immediate corollaries of Theorem 1.3 and Proposition 1.5. First, SHF($\theta$) enjoys a few invariants. The finite-dimensional (in time) analogs of these invariants were obtained in [CSZ23a, Theorem 1.2]. Let $(\tau_{z}g)(x):=g(x-z)$, with $x,z\in\mathbb{R}^{2}$, denote the shift operator on $\mathbb{R}^{2}$, and, for $r>0$, let $(\sigma^{n}_{r}f)(x):=f(x/r)$ denote the scaling operators on $\mathbb{R}^{2n}$.